the horizontal sides of a rectangle are one half the length of the side of a particular square. The vertical sides of the rectangle are 4 cm more than 3 times the side of the square. If the perimeter of the rectangle is twice the perimeter of the square, what are the demensisions of the rectangle
Found 2 solutions by ccs2011, lwsshak3:
Answer by ccs2011(207)

Let the length of the side of the square be x.
Then the dimensions of the rectangle are: 1/2x and 3x +4
Perimeter of square = 4x
Perimeter of rectangle = 1/2x + 1/2x + (3x+4) +(3x+4) = 7x + 8
Twice the perimeter of square is 8x.
Then equation is:
8x = 7x + 8
Subtract 7x on both sides
x = 8
--> 1/2(8) = 4, 3(8)+4 = 28
Therefore dimensions of rectangle are 28 by 4.

Answer by lwsshak3(11628)

the horizontal sides of a rectangle are one half the length of the side of a particular square. The vertical sides of the rectangle are 4 cm more than 3 times the side of the square. If the perimeter of the rectangle is twice the perimeter of the square, what are the demensisions of the rectangle
**
let x=horizontal side of rectangle
3(2x)+4=6x+4=vertical side of rectangle
2(6x+4)+2x=14x+8=perimeter of rectangle
..
2x=side of square
4(2x)=8x=perimeter of square
..
(14x+8)/8x=2
14x+8=16x
2x=8
x=4
ans:
dimensions of rectangle:
horizontal side: 4 cm
vertical side:6x+4=24+4=28 cm